# LEAD: Towards Learning-Based Equity-Aware Decarbonization in Ridesharing Platforms

### v1

@@ 5.1 Experimental Setup
(a) LEAD vs. baselines (b) Impact of LEVs (c) Impact of Eta ($\eta$).
@@
Figure 3: Fairness performance: (a) normalized fairness for different algorithms as a function of batch duration, (b) impact of the percentage of low emission vehicles (LEVs) in the fleet on LEAD, and (c) the impact of increasing emissions for an increase in fairness, captured using $\eta$. Here, $\eta$ specifies extra emissions that the algorithm incurs to reduce unfairness by 1km.
5.2 Impact on Per-Trip Emissions
Figure 2 shows the performance of LEAD in reducing the average emissions per trip against the baselines across various parameters.
Key results. As shown in Figure 2(a), LEAD outperforms all the baseline algorithms for all batch durations. LEAD reduces emissions per trip by at least 73.6%, 70.5%, and 58.3% when compared to CD, TORA, and LAF algorithms, respectively. LEAD does even better when the batch duration increases. While the emissions per trip for CD and TORA slightly decrease with an increase in batch duration, LEAD achieves significantly higher reductions at longer batch duration, thereby increasing the margin of improvement. Finally, LAF’s performance worsens with an increase in batch duration; emissions are 9% and 3% higher at 15 minutes than at 2 and 5 minutes, respectively.
As shown in Figure 2(b), adding low-emission vehicles (LEVs) to the ridesharing fleet reduces the average emissions per trip. Adding 10% LEVs provides up to 8.2% decrease in emissions; adding 15% LEVs reduces emissions by up to 13.5%; adding 25% LEVs achieves up to 30.9% reductions in emissions. Importantly, the lowest reductions for all LEV values occur at the longest batch duration, such as adding 10% LEVs at 15 minutes yields no added benefit.
As shown in Figure 2(c), $\eta$ does not have a significant impact on the emissions per trip. The highest increase in the emissions per trip is 5.15%, which is observed for an LEV percentage of 25%. As we increase $\eta$ to 10gCO2/km, the emissions only increase by 1.9% and 2.5% at 10% and 15% LEV percentages, respectively.
@@ 5.2 Impact on Per-Trip Emissions
Analysis of findings. We next take a deeper look at some of the results we have observed about LEAD’s impact on emissions.
  • Why does LEAD outperforms deadhead miles- and emissions-aware algorithms CD and TORA? The primary reason that LEAD outperforms CD is that CD does not take into account the unit emissions of the vehicles when assigning requests to drivers. Another reason that LEAD outperforms TORA and CD is that it assigns all the requests in each batch simultaneously, whereas TORA and CD assign the requests sequentially within each batch.
  • Why does adding LEVs yield fewer emissions reductions at longer batches? As the batch duration increases, LEAD can access a bigger set of drivers and riders, allowing it to make low-emission ride assignments, even without a high LEV percentage. The effect is more acute at smaller percentages of LEVs.
  • Why do the emissions per trip for LEAD not increase with $\eta$? The value of $\eta$ shows the grams of CO2 emitted to increase fairness by 1km. LEAD makes this tradeoff highly judiciously and only chooses the trips when improvement in fairness is possible.
Key takeaways. LEAD outperforms the state-of-the-art emission-aware ride assignment algorithm TORA by up to 80.7%. While additional reductions may be possible by optimizing batch duration and $\eta$, it will introduce tradeoffs with our other objectives of increasing driver’s fairness and reducing rider’s wait time.
@@
(a) LEAD vs. baselines (b) Impact of LEVs (c) Impact of Eta ($\eta$).
@@
Figure 4: Wait time performance: (a) wait time for different algorithms as a function of batch duration, (b) impact of the percentage of low emission vehicles (LEVs) in the fleet on LEAD performance, and (c) the impact of increasing emissions for an increase in fairness, captured using $\eta$. Here, $\eta$ specifies extra emissions that the algorithm incurs to reduce unfairness by 1km.
5.3 Impact on Driver’s Fairness
Figure 3 shows the performance of LEAD in improving fairness among drivers against the baselines across various parameters.
Key results. As shown in Figure 3(a), no other algorithm, including LEAD, beats the fairness performance of LAF. The next best fairness performance is shown by LEAD, as it reaches 90.8% of the fairness achieved by LAF at the batch duration of 5 minutes. Importantly, LEAD’s worst performance of achieving 65% of LAF’s fairness is only marginally surpassed by CD algorithm which achieves 69% of LAF’s fairness performance. TORA is the worst performing algorithm as it achieves between 37% to 49% of LAF’s fairness performance.
As shown in Figure 3(b), adding LEVs gives the greatest benefits at low batch duration, where having 10% and 15% LEVs can improve fairness by 36% and to almost 2$\times$, respectively, compared to baseline ridesharing fleet with LEAD. Increasing LEVs beyond 15% provides marginal gains in fairness across all batch durations. Finally, the gains in fairness have the lowest point at a medium batch duration of 10 minutes and increase on either side of the batch duration.
Figure 3(c) shows the effect of $\eta$ on fairness, providing us with two key observations. First, increasing $\eta$ increases fairness, with a 37% increase at 0.1g.CO2/km and 80% increase at 10g.CO2/km. Second, even a small percentage of LEVs provide a very big improvement in fairness at the higher values of $\eta$; 15% LEVs can improve fairness by 40% at $\eta$ value of 10g.CO2/km.
  • Why does fairness first increase and then decrease with increasing batch duration? As the batch duration increases, more opportunities exist to increase fairness and reduce emissions. The absolute fairness (not shown in the graph) for both LEAD and LAF increases. However, since LAF uses the benefits from batch duration only to improve fairness, while LEAD uses it for both increasing fairness and decreasing emissions, the normalized performance of LEAD with respect to LAF decreases.
  • Why does longer batch duration suppress the effect of LEVs on improving fairness? When the batch duration is short, there are fewer available drivers, limiting the algorithm to a smaller pool and reducing its ability to prioritize low-emission vehicles. However, with longer batch durations, the algorithm has a wider selection, allowing it to choose low-emission vehicles more effectively, which can lead to reduced fairness.
  • How increasing $\eta$ to improve fairness impacts emissions? As outlined in the previous section, $\eta$ has a marginal impact on emissions, with a maximum increase of 5.15%. Since increasing $\eta$ can improve fairness by almost 80%, the tradeoff seems desirable.
Key takeaways. LEAD offers a highly favorable tradeoff between improving fairness and reducing emissions. For example, at batch duration of 5 minutes, it reduces emissions by 65% while degrading fairness by less than 9%. However, using $\eta$, fairness can be made comparable with LAF for a less than 2% increase in emissions.
5.4 Impact on Rider’s Wait Time
Figure 4 shows the impact of LEAD on the wait time for riders against the baselines across various parameters. Note that CD targets reduce wait time and deadhead miles (or emission) by assigning the closest driver. TORA is explicitly wait time aware as it bounds the wait time experienced by the rider.
Key results. As shown in Figure 4(a), LEAD outperforms the baseline algorithms for all batch duration. The reduction in wait time over CD is the most significant, with 49% to 54% improvement. The performance improvement over LAF and TORA is always more than 40%. Finally, the increase in batch duration has a higher impact on our baseline algorithms, while LEAD achieves its best wait time performance at a batch duration of 5 minutes.
Figure 4(b) shows that the wait time for the riders gets worse at a higher penetration of LEVs. However, the increase in wait time, even at the worst point, is 4.16% (for 15% LEVs at 5-minute batch duration). Interestingly, increasing LEVs from 15% to 25% reduces the wait time, but by at most 1.6% and at least 0.2%.  Figure 4(c) shows that increasing $\eta$ has a small effect as well, between 2.1% to 2.8% for different penetration levels of LEVs.
Analysis of findings. We now do a deep dive into our results.
  • Why does wait time for LEAD increase after batch duration of 5 minutes? An increase in the batch duration puts a lower bound on the wait time for all the rides in the batch. However, an initial increase in batch size duration increases the possibility of a driver coming closer to the rider and becoming available. However, beyond a certain batch duration, the effect of increasing the lower-bound dominates, and wait time goes up.
  • Why having more LEVs is bad for wait time? How does it impact the usefulness of LEAD? As LEAD targets reducing emissions, it will prioritize EVs even if they are farther away or will only be available after finishing a ride, thus increasing the wait time. While the increase in wait time is not desirable, the increase is not significant, and the addition of LEVs provides significant emission reduction and fairness benefits.
Key takeaways. LEAD outperforms CD by considering the future availability of the drivers, same as TORA. However, its explicit focus on reducing deadhead miles reduces wait time beyond TORA. Importantly, LEAD’ wait time is less sensitive to batch duration and $\eta$, allowing for reducing emissions and increasing fairness.
5.5 Additional Performance Benchmarking
In this section, we benchmark LEAD and other baselines using the additional metrics that evaluate system-level performance. The results are shown in Table 2 and Table 3.
LEAD outperforms baseline algorithms in improving utility by at least 19% and reducing deadhead-to-trip ratio by 15.2% over the second-best algorithm LAF. However, LAF rejects many ride requests, as outlined in the original paper, and accepts only 93.3% requests compared to 98.3% ride acceptance rate for LEAD.
Importantly, LEAD also solves TORA’s problem of assigning a disproportionate number of rides to LEVs by allocating rides to LEVs in proportion to their penetration in the fleet.
Table 2: Performance comparison of LEAD against baselines for $\eta=5$ g.CO2/km and batch duration of $5$ minutes.
Algorithm Avg. utilities (km) Avg. deadhead Percentage of
to trip ratio matched requests
CD 763.14 0.66 $97.4\%$
TORA [32] 820.4 0.62 $97.3\%$
LAF [35] 890.56 0.59 $93.3\%$
LEAD (ours) 1065.09 0.50 $98.3\%$
Table 3: Fraction of rides assigned to LEVs at various LEV %ages for $\eta$=5 g.CO2/km and batch duration of $5$ minutes.
Algorithm Original 10% LEV 15% LEV 25% LEV
TORA [32] 13.58 20.41 26.07 35.05
LEAD (ours) 6.9 11.60 16.61 26.16
6 Discussion and Future Works
@@
In this paper, we propose LEAD, a Learning-based Equity-Aware Decarbonization approach for ridesharing platforms. LEAD minimizes emissions while ensuring fair distribution of drivers utility. LEAD leverages reinforcement learning to optimize rider-driver matches based on the expected future utility for drivers and projected carbon emissions for the platform without increasing rider waiting times. Extensive experiments using a public real-world ridesharing dataset demonstrate that LEAD outperforms state-of-the-art baselines in terms of reduction in emissions and wait time while providing a desirable tradeoff with fairness. In the future, we will explore fairness from a rider’s perspective and short-term fluctuations in driver utility, in addition to the long-term effect this paper considers.
7 Acknowledgements

### v2

@@ 5.3. Impact on Driver’s Fairness
(a) LEAD vs. baselines (b) Impact of LEVs (c) Impact of parameter $\eta$
@@
Figure 3. Fairness performance: (a) normalized fairness for different algorithms as a function of batch duration, (b) impact of the percentage of low emission vehicles (LEVs) in the fleet on LEAD for different batch durations and $\eta$ = 5, and (c) the impact of increasing emissions for an increase in fairness, captured using $\eta$, for a batch duration of 5 minutes. Here, $\eta$ specifies extra emissions that the algorithm incurs to reduce unfairness by 1km.
  • What causes the significant increase in fairness when the proportion of LEVs rises from 15% to 25% compared to baseline LEAD?
    Fairness increases with the addition of 25% LEVs compared to the baseline LEAD due to a availability of more number of LEVs. When the proportion of LEVs is 10% or 15%, the system prioritizes assigning most ride requests to LEVs to reduce emissions, which leads to unfairness. However, with 25% LEVs, the system achieves a better balance in ride assignments between LEVs and HEVs, resulting in significantly improved fairness.
  • How increasing $\eta$ to improve fairness impacts emissions?
    As outlined in the previous section, changing $\eta$ between $0.1$ to $10$ gCO2/km has only a marginal impact on emissions, with a maximum increase of 0.8%. Since increasing $\eta$ can lead to significant improvements in fairness by almost 70%, the tradeoff seems desirable.
Key takeaways. LEAD offers a highly favorable tradeoff between improving fairness and reducing emissions. For example, when compared to LAF, at batch duration of 10 minutes, it reduces emissions by 56.6% while degrading fairness by 29%.
5.4. Impact on Rider’s Wait Time
Figure 4 shows the impact of LEAD on the wait time for riders against the baselines across various parameters. Note that CD targets reduce wait time and deadhead miles (or emission) by prioritizing the assignment of the closest available driver. On the other hand, TORA is designed with explicit consideration for wait times. It incorporates mechanisms to bound the wait time experienced by the rider.
Key results. As shown in Figure 4 (a), LEAD outperforms LAF in terms of reducing rider wait times across all batch durations. Compared to LAF, LEAD shows significant performance gains, with improvements consistently exceeding 32%.
Figure 4 (b) shows that the wait time for the riders gets worse at a lower penetration of LEVs. For example, the increase in wait time, even at the worst point, is 4.2% for 10% LEVs at 10-minute batch duration. Interestingly, increasing LEVs from 15% to 25% reduces the wait time, but by at most 7.5% and at least 4.5%.
Figure 4 (c) shows that increasing $\eta$ has a small effect on the rider’s waiting time as well, the increase in waiting time is between 2.2% and 3.1% for 15% LEVs, and the decrease in waiting time is between 6.5% and 7.8% for 25% LEVs.
@@ 5.4. Impact on Rider’s Wait Time
Analysis of findings. We next delve deeper into some results observed about LEAD’s impact on wait time.
  • Why does wait time for LEAD increase as the batch duration increases?
    Longer batch durations delay the next assignment, causing riders to wait more, even if closer drivers become available in the meantime. This leads to an overall increase in average wait time for riders.
  • Why having more LEVs is good for wait time? How does it impact the usefulness of LEAD?
    As LEAD focuses on reducing emissions, it prioritizes assigning rides to LEVs, even if they are farther from the passenger, currently serving another ride request, or will only become available later. When the percentage of LEVs is low, finding nearby LEVs is challenging, leading to increased passenger wait times. However, as the proportion of LEVs increases, such as to 25%, the algorithm has access to a larger pool of LEVs, making it easier to locate available LEVs closer to passengers, thereby reducing wait times and enhancing the algorithm’s effectiveness.
Key takeaways. LEAD outperforms LAF by considering the future availability of the drivers. However, what sets LEAD apart is its reduced sensitivity to batch duration and $\eta$ with respect to wait times. This allows LEAD to achieve lower emissions and improved fairness.
@@
(a) LEAD vs. baselines (b) Impact of LEVs (c) Impact of parameter $\eta$
@@
Figure 4. Wait time performance: (a) wait time for different algorithms as a function of batch duration, (b) impact of the percentage of low emission vehicles (LEVs) in the fleet on LEAD performance for different batch durations and $\eta$ = 5, and (c) the impact of increasing emissions for an increase in fairness, captured using $\eta$, for a batch duration of 5 minutes. Here, $\eta$ specifies extra emissions that the algorithm incurs to reduce unfairness by 1km.
6. Conclusion
@@
In this paper, we propose LEAD, a Learning-based Equity-Aware Decarbonization approach for ridesharing platforms. The primary objective of LEAD is to minimize carbon emissions while ensuring a fair distribution of drivers utility. By leveraging reinforcement learning, LEAD optimizes the rider-driver matching process based on the expected future utility the driver could achieve and the projected carbon emissions for the platform without an increase in rider waiting times. Through extensive experiments using a publicly available real-world ridesharing dataset and a variety of problem and algorithm parameters, we demonstrate that LEAD outperforms state-of-the-art baselines in terms of reduction in emissions and wait time while ensuring a desirable level of fairness among drivers with varying vehicle emissions. In the future, we will focus on expanding our fairness approach to include the rider’s perspective, ensuring equitable ridesharing service across different rider demographics and locations. Additionally, we plan to investigate short-term fluctuations in driver utility, complementing the long-term impacts explored in this paper, to better understand the dynamic nature of driver income and utility within ridesharing platforms.
7. Acknowledgements